# On the distance to the black hole X-ray binary Swift J1727.8--1613

### v1

@@ Abstract
We review the existing distance estimates to the black hole X-ray binary Swift J1727.8--1613 (catalog ) with a discussion of the accuracies and caveats of the associated methodologies. As part of this, we present new line-of-sight H i absorption spectra captured using the MeerKAT radio telescope. We estimate a maximum radial velocity with respect to the local standard of rest of $24.8\pm 2.8$ km s-1, which is significantly lower than that found towards an extragalactic reference source. Given the location of Swift J1727.8--1613 at Galactic longitude and latitude $(l,b)\approx(8.6^{\circ},10.3^{\circ})$, we explore the feasibility of the H i absorption method. From this we derive a near kinematic distance of $d_{\rm{near}}=3.6\pm 0.3\,({stat})\pm 2.3\,({sys})$ kpc as lower bound for the distance to Swift J1727.8--1613. We compare our results with those derived from different distance determination methods including the use of colour excess or reddening along the line of sight, which we constrain to $E(B\!-\!V)=0.37\pm 0.01\,({stat})\pm 0.025\,({sys})$ using near-UV spectra. By combining this with donor star magnitudes reported by Mata Sánchez et al. (2025), we suggest an increased distance of $5.5^{+1.4}_{-1.1}$ kpc, which would imply a natal kick velocity of $190\pm 30$ km s-1.
@@
facilities: MeerKAT, HST/STISsoftware: Astropy (Astropy Collaboration et al., 2022), CASA (CASA Team et al., 2022), CARTA (Comrie et al., 2021), dustmaps (Green, 2018), Jupyter (Kluyver et al., 2016), kd (Wenger, 2018), Matplotlib (Hunter, 2007), Miriad (Sault et al., 1995), NumPy (Harris et al., 2020), PyMC (Abril-Pla et al., 2023), SciPy (Virtanen et al., 2020), stistools (Hack et al., 2018).
[…]
revtex4-1.clsrevtex4-2.cls
@@ 1 Introduction
Distance measurement is important for all objects. In studying Galactic low-mass X-ray binaries (XRBs), accurate distances allow for better estimation of parameters such as the peak Eddington luminosity ($L_{\rm{Edd}}$) fraction (ELF) and jet parameters including physical size scales, inclination angles, and speeds.
Reliably measuring the distance to newly discovered XRBs can be challenging. For instance, as the accretion processes of low-mass XRBs evolve, the system’s luminosity can fluctuate, often in and out of detectable levels. This inconsistent detectability can preclude the measurements required for accurate distance determination.
@@ 1.1 Distance methods
Distance determination techniques for XRBs can be broadly summarised as methods that use either astrometry or kinematics, observations of the donor star or X-ray source, or propagation effects.
@@ 1.1.1 Astrometry and kinematics
High-significance XRB parallax measurements with Gaia (Gandhi et al., 2019; Atri et al., 2019; Arnason et al., 2021) or very long baseline interferometry (VLBI) at radio wavelengths (e.g.; Miller-Jones et al., 2009; Reid et al., 2011; Atri et al., 2020; Miller-Jones et al., 2021; Reid & Miller-Jones, 2023) are the gold standard for measuring distances. However, radio parallaxes can be impeded by line-of-sight scatter broadening for XRBs located in the Galactic Plane (GP). Given the typical kpc distances of XRBs, sub-milliarcsecond precision is required (Tetarenko et al., 2016).
Alternatively, kinematic distance methodologies use measured proper motions and velocities predicted using the Galactic rotation model to infer the most likely distance. In some circumstances, these have been shown to have similar if not better accuracy than parallax (Reid, 2022) and useful for XRBs (e.g.; Dhawan et al., 2007; Reid & Miller-Jones, 2023).
1.1.2 Stellar and X-ray observations
One can use optical spectroscopy of the XRB donor star to estimate the distance (e.g.; Dubus et al., 2001; Jonker & Nelemans, 2004; Charles et al., 2019). With measured values for the donor star’s absolute and apparent magnitudes and the line-of-sight extinction along the line of sight, one can then use the distance modulus (e.g.; Mata Sánchez et al., 2024, 2025)
@@ 1.1.2 Stellar and X-ray observations
Obscuration of optical light is pronounced for targets residing in the GP due to increased interstellar dust, making donor stars difficult to identify, and introducing additional uncertainty to the distance modulus equation. However, for targets outside the GP, extinction can be harder to estimate accurately. We use and discuss this method in Section 4.2.
@@
X-ray luminosities of XRB outbursts during soft-to-hard and hard-to-intermediate state transitions have been observed to occur at somewhat consistent ELFs, albeit with factor of $\sim\!3$ scatter in these measurements (Kalemci et al., 2013; Tetarenko et al., 2016; Vahdat Motlagh et al., 2019). X-ray studies during these transitions allow the comparison of measured and expected intrinsic luminosities, enabling the estimation of distance (e.g.; Abdulghani et al., 2024).
1.1.3 Propagation
By analysing the time delays and intensities of X-rays that had been produced by flares and subsequently scattered off intervening interstellar dust clouds, one can reconstruct the XRB outburst light curve and determine the distance to the source (e.g.; Heinz et al., 2015; Beardmore et al., 2016; Lamer et al., 2021).
@@ 1.1.3 Propagation
X-ray absorption features in observed spectra can be used to measure the hydrogen column density, $N_{\rm{H}}$. When coupled with hydrogen distribution models one can infer the distance to the source. We discuss this further in Section 4.2.
The relation between the colour excess or reddening, $E(B\!-\!V)$, and extinction can also be used to inform distance modulus calculations (e.g.; Schlafly & Finkbeiner, 2011). The inverse relation between $E(B\!-\!V)$ and the distance therefore allows constraints on one to constrain the other. $E(B\!-\!V)$ can be estimated using various relations between it and interstellar absorption lines (e.g.; Munari & Zwitter, 1997; Wallerstein et al., 2007), or from near-UV observations, as used in Sections 2.2 and 3.2. $E(B\!-\!V)$ has also been determined from 2-dimensional (2D; e.g.; Schlegel et al., 1998; Chiang, 2023) and 3-dimensional (3D; e.g.; Green et al., 2019; Edenhofer et al., 2024) Galactic dust maps.
@@
Lastly, line-of-sight H i absorption has long been used as an XRB distance estimator (e.g.; Dickey, 1983; Lockman et al., 2007; Chauhan et al., 2019, 2021). We have also applied this method in Sections 2.1 and 3.1.
MJD Observation Observation Exposure Centre Total Channel J1727
@@
start date start time time frequency bandwidth width peak flux density
(dd-mm-yyyy) (hh:mm:ss) (mm:ss) (MHz) (MHz) (kHz) (mJy)
60183 27-08-2023 15:27:59.6 14:55.6 1283.9869 856 26.123 $49.7\pm 0.2$
60193 06-09-2023 15:06:32.1 14:56.9 1419.9984 107 3.265 $97.5\pm 0.4$
60231 14-10-2023 12:40:20.2 14:55.6 1283.9869 856 26.123 $836.6\pm 2.3$
60233 16-10-2023 15:50:13.8 14:56.9 1419.9984 107 3.265 $104.3\pm 0.4$
@@
1.1.4 H I absorption
H i absorption occurs when clouds of neutral hydrogen along the line of sight absorb the broadband continuum emission produced by the target at the H i frequency in their rest frame. These clouds move with different velocities relative to us along the line of sight, due to the rotation of the Milky Way, as well as other effects such as non-circular streaming motions that we assume to be minimal. The more clouds that are intersected by the line of sight, the more H i absorption features that are imprinted at different frequencies on the observed radio spectrum. One benefit of H i absorption over parallax is that the required data can be gathered within a much shorter timeframe; a single observation can suffice should the observed source be particularly bright.
Using the Milky Way rotation curve, the Doppler-shifted frequencies can be converted into local standard of rest (LSR) velocities. The maximum velocity occurs at the tangent point, where the rotational velocity is entirely along the line of sight. Identical velocities are seen either side of this maximum, giving rise to an ambiguity in mapping observed absorption velocities to distances within the solar circle. A maximum observed velocity that is less than the tangent point velocity could correspond to a near kinematic distance before the tangent point, or a far kinematic distance beyond the tangent point (e.g.; Wenger et al., 2018, Figure 4).
To resolve this kinematic distance ambiguity, one must observe the target but also at least one extragalactic reference source close enough to the target in the sky such that any differences in the anticipated H i distributions along the lines of sight are minimised. The emission from the reference source will have passed through all Galactic H i clouds along the line of sight, with clouds outside the solar circle imprinting absorption velocities of the opposite sign. Any absorption present in the reference spectrum but absent in the target spectrum then allows us to place an upper limit on the distance. Further to this, observations of two-sided relativistic jets can be used to give a maximum distance to a source.
1.2 Swift J1727.8--1613
Swift J1727.8--1613 (J1727), located at $(l,b)=(8.641502^{\circ},10.254899^{\circ})$, was first detected as an X-ray transient on 24 August 2023 (Negoro et al., 2023), after which bright radio emission was observed within a couple of days (Miller-Jones et al., 2023b), which continued to brighten through early September (Bright et al., 2023). Analysis of observations in late August and early September revealed a bright core and a large two-sided, asymmetrical jet (Wood et al., 2024). Radio monitoring in early October suggested radio quenching and subsequent flaring (Miller-Jones et al., 2023a).
This event was deemed a low-mass XRB outburst (Castro-Tirado et al., 2023), and its high radio brightness made J1727 a suitable target for H i absorption measurements. Since then, the compact object has been dynamically confirmed to be a black hole (BH) (Mata Sánchez et al., 2025, MS24b hereafter).
Abdulghani et al. (2024) provided an initial distance estimate of $1.52^{+0.85}_{-0.61}$ kpc from the ELF method mentioned in Section 1.1.2. The authors concede that this particular distance estimate may be underestimated by up to $\sim\!70\%$. Mata Sánchez et al. (2024, MS24a hereafter) suggested a distance of $2.7\pm 0.3$ kpc. MS24b classified the donor star type as K3/4V and updated the distance to $3.4\pm 0.3$ kpc.
The Gaia optical counterpart for Swift J1727.8--1613 currently has proper motion but no parallax. A VLBI radio parallax will not be possible, as the XRB has already returned to the quiescent state.
In Section 2 we detail the methods used and data obtained. In Section 3 we present our H i absorption and near-UV spectra. In Section 4 we discuss the interpretation of the H i absorption results, their use in estimating the kinematic distance, and the associated caveats. We extend this discussion to compare our result with others published to date, augment this using our near-UV results, and explore the implications of the resulting distance estimate. In Section 5 we present our best estimate of the distance to J1727.

### v2

# On the distance to the black hole X-ray binary Swift J1727.8-1613

@@ Abstract
We review the existing distance estimates to the black hole X-ray binary Swift J1727.8-1613 (catalog ), present new radio and near-UV spectra to update the distance constraints, and discuss the accuracies and caveats of the associated methodologies.
We use line-of-sight H i absorption spectra captured using the MeerKAT radio telescope to estimate a maximum radial velocity with respect to the local standard of rest of $24.8\pm 2.8$ km s-1 for Swift J1727.8-1613, which is significantly lower than that of a nearby extragalactic reference source. From this we derive a near kinematic distance of $d_{\rm{near}}=3.6\pm 0.3\,({stat})\pm 2.3\,({sys})$ kpc as a lower bound after accounting for additional uncertainties given its Galactic longitude and latitude, $(l,b)\approx(8.6^{\circ},10.3^{\circ})$.
Near-UV spectra from the Hubble Space Telescope’s Space Telescope Imaging Spectrograph allows us to constrain the line-of-sight colour excess to $E(B\!-\!V)=0.37\pm 0.01\,({stat})\pm 0.025\,({sys})$. We then implement this in Monte Carlo simulations and present a distance to Swift J1727.8-1613 of $5.5^{+1.4}_{-1.1}$ kpc, under the assumption that the donor star is an unevolved, main sequence K4($\pm$1)V star. This distance implies a natal kick velocity of $190\,\pm\,30$ km s-1 and therefore an asymmetrical supernova explosion within the Galactic disk as the expected birth mechanism.
A lower distance is implied if the donor star has instead lost significant mass during the binary evolution. Hence, more accurate measurements of the binary inclination angle or donor star rotational broadening from future observations would help to better constrain the distance.
@@
facilities: MeerKAT, HST/STISsoftware: ArviZ (Kumar et al., 2019), Astropy (Astropy Collaboration et al., 2022), CASA (CASA Team et al., 2022), CARTA (Comrie et al., 2024), dustmaps (Green, 2018), Jupyter (Kluyver et al., 2016), kd (Wenger, 2018), Matplotlib (Hunter, 2007), Miriad (Sault et al., 1995), NumPy (Harris et al., 2020), PreliZ (Icazatti et al., 2023), PyMC (Abril-Pla et al., 2023), SciPy (Virtanen et al., 2020), stistools (Hack et al., 2018).
@@ 1 Introduction
Distance is an important parameter in the study of all astrophysical objects. For Galactic low-mass X-ray binaries (XRBs), accurate distances allow for better estimation of other parameters, such as the peak Eddington luminosity ($L_{\rm{Edd}}$) fraction (ELF) and jet parameters including physical size scales, inclination angles, and speeds.
Reliably measuring the distance to newly discovered XRBs can be challenging. For instance, across the different phases of an outburst, the system’s luminosity can fluctuate, often in and out of detectable levels. This inconsistent detectability can preclude the measurements required for accurate distance determination.
@@ 1.1 Distance methods
Distance determination techniques applicable to XRBs can be broadly categorised into three groups: (1) astrometric or kinematic methods; (2) approaches that use observations of the donor star, X-ray source, or jets; or (3) techniques based on propagation effects. We summarise these in Sections 1.1.11.1.3. We then discuss in more detail the two techniques that we employ that exploit the relationships between the distance and H i absorption in radio observations (Section 1.1.4), and the colour excess or reddening, E(BV)E(B\!-\!V), as measured using near-UV observations (Section 1.1.5).
@@ 1.1.1 Astrometry and kinematics
High-significance XRB parallax measurements with Gaia (Gaia Collaboration et al., 2016; Gandhi et al., 2019; Atri et al., 2019; Arnason et al., 2021) or very long baseline interferometry (VLBI) at radio wavelengths (e.g.; Miller-Jones et al., 2009; Reid et al., 2011, 2014a; Atri et al., 2020; Miller-Jones et al., 2021; Reid & Miller-Jones, 2023) are the gold standard for measuring distances to Galactic XRBs. However, radio parallaxes can be impeded by line-of-sight scatter broadening for XRBs located in the Galactic Plane (GP). In fact, extinction in the GP and the faintness of quiescent XRBs precludes Gaia distances in many cases. Given the typical kiloparsec (kpc) distances of XRBs, sub-milliarcsecond precision is required (Tetarenko et al., 2016). These methods also require observations that span an extended timeframe, which is not always possible for XRB outbursts.
Alternatively, kinematic distance methodologies use measured proper motions and velocities predicted using the Galactic rotation model to infer the most likely distance. Despite relying on the assumption of low peculiar velocities relative to the local standard of rest (LSR), these can provide reliable distances in some circumstances (Reid, 2022) and can be useful for XRBs (e.g.; Dhawan et al., 2007; Reid & Miller-Jones, 2023).
1.1.2 Stellar, X-ray, and jet observations
One can use optical spectroscopy of the XRB donor star to estimate the distance (e.g.; Dubus et al., 2001; Jonker & Nelemans, 2004; Charles et al., 2019). With measured values for the donor star’s absolute and apparent magnitudes and the extinction along the line of sight, one can use the distance modulus (e.g.; Mata Sánchez et al., 2024, 2025) to infer the distance, via
@@ 1.1.2 Stellar, X-ray, and jet observations
Obscuration of optical light is pronounced for targets residing in the GP due to increased interstellar dust, making donor stars difficult to identify, and introducing additional uncertainty to the distance modulus equation. However, for targets outside the GP, extinction can be harder to estimate accurately.
@@
X-ray luminosities of XRB outbursts during soft-to-hard and hard-to-intermediate state transitions have been observed to occur at somewhat consistent ELFs, albeit with factor of $\sim\!3$ scatter in these measurements (Kalemci et al., 2013; Tetarenko et al., 2016; Vahdat Motlagh et al., 2019). X-ray studies during these transitions allow one to compare the measured and expected intrinsic luminosities and thereby estimate the distance (e.g.; Abdulghani et al., 2024).
Further X-ray methods exist, including the combination of X-ray spectroscopy and the distance dependency of accretion disk spectral fits, which Hynes et al. (2002) applied to constrain the distance to XTE J1859+226. Additionally, Powell et al. (2007) used the time-scale of outburst decay X-ray light curves to estimate the absolute luminosity at a characteristic time and therefore provide a measure of the distance.
In the radio band, the proper motions of two-sided jets can be combined to place an upper limit on the source distance (e.g., Mirabel & Rodríguez, 1994).
1.1.3 Propagation and the Interstellar Medium
X-rays produced by XRB flares will propagate outwards and may subsequently scatter off intervening interstellar dust clouds. Provided the distances to these dust clouds can be determined, one can combine this information with analysis of the time delays and intensities of these expanding X-ray dust scattering rings to measure the distance to the source (e.g.; Heinz et al., 2015; Beardmore et al., 2016; Lamer et al., 2021).
@@ 1.1.3 Propagation and the Interstellar Medium
X-ray absorption features in observed spectra can be used to measure the hydrogen column density, $N_{\rm H}$. When coupled with hydrogen distribution models one can infer the distance to the source.
The relation between E(BV)E(B\!-\!V) and the aforementioned extinction along the line of sight can be used to inform distance modulus calculations (e.g.; Schlafly & Finkbeiner, 2011). The inverse relation between E(BV)E(B\!-\!V) and the distance therefore allows constraints on one to constrain the other. We examine and implement this method in Sections 1.1.5 and 4.2 respectively.
@@
Lastly, line-of-sight H i absorption has long been used as an XRB distance estimator (e.g.; Dickey, 1983; Lockman et al., 2007; Chauhan et al., 2019, 2021). We explore this method further in Section 1.1.4 below and apply it in Sections 2.1, 3.1, and 4.1.
1.1.4 H I absorption
Our first distance method uses H i absorption, which can be observed when clouds of neutral hydrogen along the line of sight absorb the broadband continuum emission produced by the target at the H i frequency in their rest frame. These clouds move with different velocities relative to us along the line of sight, due to the rotation of the Milky Way, as well as other effects such as non-circular streaming motions that we assume to be minimal. The more clouds that are intersected by the line of sight, the more H i absorption features that are imprinted at different frequencies on the observed radio spectrum. One benefit of H i absorption over parallax is that the required data can be gathered within a much shorter timeframe; a single observation can suffice should the observed source be particularly bright.
The Doppler-shifted frequencies can be converted into LSR velocities and compared to the Milky Way rotation curve. The maximum velocity occurs at the tangent point, where the rotational velocity is entirely along the line of sight. Identical velocities are seen either side of this maximum, giving rise to an ambiguity in mapping observed absorption velocities to distances within the solar circle. A maximum observed velocity that is less than the tangent point velocity could correspond to a near kinematic distance before the tangent point, or a far kinematic distance beyond the tangent point (e.g.; Wenger et al., 2018, Figure 4).
To resolve this kinematic distance ambiguity, one must observe the target but also at least one extragalactic reference source close enough to the target in the sky such that any differences in the anticipated H i distributions along the lines of sight are minimised. The emission from the reference source will have passed through all Galactic H i clouds along the line of sight, with clouds outside the solar circle imprinting absorption velocities of the opposite sign. Any absorption present in the reference spectrum but absent in the target spectrum then allows us to place an upper limit on the distance.
1.1.5 E(BV)E(B\!-\!V)
Our second distance method relies on the reddening caused by interstellar dust preferentially scattering shorter wavelengths. This can be determined by subtracting the observed difference between blue and visible magnitudes, $B$ and $V$, to quantify the reddening along the line of sight.
E(BV)E(B\!-\!V) can be calculated using various relations between it and interstellar absorption lines (e.g.; Munari & Zwitter, 1997; Wallerstein et al., 2007), or $N_{\rm H}$ (e.g.; Mata Sánchez et al., 2025). It can be estimated from Galactic dust maps, both 2-dimensional (2D; e.g.; Schlegel et al., 1998; Chiang, 2023) and 3-dimensional (3D; e.g.; Green et al., 2019; Edenhofer et al., 2024). Near-UV spectra can also be used as implemented in Sections 2.2 and 3.2.
With a value for E(BV)E(B\!-\!V) one can derive the extinction. For example, it is common to use $R_{V}=3.1$ as a Galactic average with
$A_{V}=R_{V}\,E(B\!-\!V)$ (2)
to convert reddening to the total extinction along the line of sight or vice versa (Savage & Mathis, 1979; Fitzpatrick, 2004). When combined with measurements of absolute and apparent magnitudes in Equation 1, the distance can then be calculated.
1.2 Swift J1727.8-1613
Swift J1727.8-1613 (J1727), located at $(l,b)=(8.641502^{\circ},10.254899^{\circ})$, was first detected as an X-ray transient on 24 August 2023 (Negoro et al., 2023). Bright radio emission was observed within a couple of days (Miller-Jones et al., 2023b), which continued to brighten through early September (Bright et al., 2023). Analysis of observations in late August and early September revealed a bright core and a large two-sided, asymmetrical jet (Wood et al., 2024). Radio monitoring in early October suggested radio quenching and subsequent flaring (Miller-Jones et al., 2023a).
This event was deemed a low-mass XRB outburst (Castro-Tirado et al., 2023), and its high radio brightness made J1727 a suitable target for H i absorption measurements. Since then, the compact object has been dynamically confirmed to be a black hole (BH) (Mata Sánchez et al., 2025, MS25 hereafter).
Further studies of the outburst revealed that it produced relativistic jets that are the largest resolved jets in an XRB to date (Wood et al., 2024). The ejection of transient jets in J1727 has also been shown to have occurred simultaneously with a bright X-ray flare and a sudden change in the X-ray properties of the accretion inflow (Wood et al., 2025).
The Gaia optical counterpart for Swift J1727.8-1613 currently has proper motion but no parallax. A VLBI radio parallax will not be possible with the observations taken to date, as the XRB has already returned to the quiescent state.
1.2.1 Current distance estimates
Abdulghani et al. (2024) estimated a distance of $1.52^{+0.85}_{-0.61}$ kpc from a Bayesian approach of soft-state X-ray modelling. This appeared to align with Veledina et al. (2023) who used X-ray flux scaling arguments to provide an early estimate of approximately $1.5$ kpc. However, Abdulghani et al. (2024) concede that their distance estimate may be underestimated by up to $\sim\!70\%$, given that only soft-state data and no state-transition information was used.
Mata Sánchez et al. (2024, MS24 hereafter) used donor star magnitudes in conjunction with various relations in the literature to derive values for the parameters in Equation 1. These included the relation between the interstellar Ca ii doublet (H and K) and the distance to early-type stars per Megier et al. (2009). This was calibrated using objects within a few hundred pc from the GP (up to 450 pc), slightly below but still consistent with their final inferred height for J1727. The authors also used the relations between the equivalent widths of the interstellar line K i 7699 Å and diffuse interstellar band at 8621 Å and E(BV)E(B\!-\!V) per Munari & Zwitter (1997) and Wallerstein et al. (2007) respectively. However, these relations resulted in particularly large values for E(BV)E(B\!-\!V) at $0.8\pm 0.3$ and $0.9\pm 0.3$ respectively. Lastly, the relation between hydrogen column density $N_{\rm H}$ and $V$-band extinction $A_{V}$ per Güver & Özel (2009) was combined with Equation 2 to infer $E(B\!-\!V)=0.47\pm 0.13$. These values of E(BV)E(B\!-\!V) cover a wide range and some have large associated uncertainties, which propagate to the distance constraints. We discuss this further in Section 4.2.1 using our near-UV results and Galactic dust maps.
Following the above, MS24 calculated the weighted mean of the resulting distances to be $d=2.7\pm 0.3$ kpc. MS25 then directly measured the orbital period, $P_{\rm orb}$, and reported the best fitting spectral type template of K4($\pm$1)V for a donor star that is partially veiled by the accretion disk. Using this, they revised the absolute $r$-band magnitude to $M_{r}=6.6\pm 0.5$ mag. They also measured an apparent $r$-band magnitude of $m_{r}=21.13\pm 0.05$ mag, and presented an updated consolidated weighted mean distance of $d=3.4\pm 0.3$ kpc.
1.2.2 Contents
MJD Observation Observation Exposure L-band Centre Total Channel J1727 peak
@@ 1.2.2 Contents
start date start time time mode frequency bandwidth width flux density
($d$-$m$-yyyy) ($h$:$m$:$s$) ($m$:$s$) (MHz) (MHz) (kHz) (mJy)
60183 27-08-2023 15:27:59.6 14:55.6 Standard 1283.9869 856 26.123 $49.7\pm 0.2$
60193 06-09-2023 15:06:32.1 14:56.9 Zoom 1419.9984 107 3.265 $97.5\pm 0.4$
60231 14-10-2023 12:40:20.2 14:55.6 Standard 1283.9869 856 26.123 $836.6\pm 2.3$
60233 16-10-2023 15:50:13.8 14:56.9 Zoom 1419.9984 107 3.265 $104.3\pm 0.4$
@@
In Section 2 we detail the methods used and data obtained. In Section 3 we present our H i absorption and near-UV spectra. In Section 4 we discuss the interpretation of our results in constraining the distance to J1727, along with various caveats and implications for natal kick velocities and Eddington luminosity fractions. In Section 5 we present our suggested distance to J1727.